# Simon Yu1  Gang Li2  Weiyan Shi†1  Peng Qi†2

### v1

@@ 4.2 Analysis
To investigate how skills are induced during training, we use our newly proposed metrics to track the model’s learning dynamics. This analysis builds on previous work that sought to understand these dynamics (Shah et al., 2025). For our study, we saved 20 model snapshots at regular intervals while training on the Mind2Web Cross-task and WebArena Shopping benchmarks. We then evaluated our metrics on each snapshot to observe how skills emerge over time. The following plots illustrate these trends.
@@
To validate the fundamental hypothesis that skill learning improves task efficiency, we analyze the relationship between Skill Reusability and the number of steps required for task completion. The results, presented in Figure 5, demonstrate a clear inverse correlation across all three methods: as skill reusability increases from 0% to over 20%, the average number of steps decreases substantially from approximately 6.1 to 3.3-4.4 steps.
@@
PolySkill achieves the highest Skill Reusability (reaching 20.4% by task 180), while maintaining competitive step reduction. Notably, ASI shows the most dramatic step reduction despite lower utilization rates, suggesting efficient skill application. This analysis confirms that learned skills directly translate to improved task efficiency, with higher utilization rates consistently leading to more streamlined task execution. The strong correlation validates our core premise that skill induction and reuse are key drivers of agent performance improvement in complex web navigation tasks.
Figure 6: A continual learning experiment showing PolySkill can prevent catastrophic forgetting. The experiment consists of two phases: initial skill library induced on the in-domain WebArena Shopping benchmark; followed by continual learning on cross-webiste: Amazon then Target. The orange and red lines (right y-axis) show that PolySkill learns the new websites more effectively than the ASI baseline. The blue lines (left y-axis) track performance on the original WA results throughout the experiment. Shaded regions represent the standard error across three runs.
@@
Case Study: Continual Learning To simulate how skills would be used and evolved in real-world scenarios, we investigate the agent’s continual learning capabilities. The experiment starts with an agent whose skill library was learned on the WebArena Shopping tasks. The agent then continues to perform online updates via skill induction on new websites for Amazon and then Target, respectively (task details in Section C.4). To reduce the potential noise, we performed the experiments 3 times to reduce the potential variance across runs. In this setting, we are both interested in the positive transfer from the existing skill library in a similar domain. Also, we measure the WA Shopping performance after it has adapted to the new website. This allows us to study potential catastrophic forgetting, where learning new knowledge can harm existing skills, a well-known challenge in continual learning (van de Ven et al., 2025).
@@ 5 Conclusion
Looking ahead, this framework opens several avenues for research, including handling highly dynamic websites, enabling skill sharing between agents, and incorporating human feedback. More broadly, the principle of polymorphism is not limited to web agents. It offers a powerful template for any agent that must operate in diverse environments sharing similar underlying structures (Xu et al., 2025): from robotics, where skills must generalize across different physical settings (Cheng et al., 2025; Liu et al., 2025a), to tool use, where software interfaces constantly evolve (Fei et al., 2025; Qiu et al., 2025). This work, therefore, provides a concrete step toward building more robust and adaptive agents capable of learning from experience (Silver & Sutton, 2025) in diverse environments.
6 Limitations and Future Work
6.1 Limitations
While PolySkill demonstrates significant improvements in skill generalization, several challenges remain that merit discussion.
Robustness to Dynamic Web Environments
Our approach handles variations in static website layouts effectively, but faces challenges with highly dynamic web applications. On sites where page structures change frequently, concrete skill implementations can quickly become obsolete. This requires costly re-validation and re-induction cycles to maintain performance, limiting the framework’s practical applicability in fast-moving web environments.
Quality of Abstract Class Initialization
The effectiveness of our framework hinges on successfully inducing high-quality abstract classes during the initialization phase. This process can be sensitive and often requires multiple successful trajectories across similar websites to produce robust, truly general interfaces. When the initial abstraction captures superficial similarities rather than semantic ones, all subsequent concrete implementations inherit these limitations, creating a systematic weakness that propagates throughout the skill hierarchy.
Coverage of the Long Tail
The polymorphic structure excels in well-defined domains with clear shared patterns, such as e-commerce or developer platforms. However, its utility diminishes for rare ”long-tail” websites that don’t fit existing categories, or for complex sites that blend functionalities from multiple domains. For example, a website that combines social networking, e-commerce, and content creation might not align well with any single abstract class, limiting the framework’s ability to transfer skills effectively.
@@ 6.2 Future Work
These limitations point to several promising research directions that could significantly enhance the framework’s capabilities.
Adaptive Skill Repair Mechanisms
To address the challenge of dynamic web environments, future work should explore automatic skill repair mechanisms. Rather than requiring full re-induction when layouts change, the system could use differential analysis to identify minimal modifications needed to restore functionality. This could involve comparing broken implementations against working ones to pinpoint the source of failure, or leveraging visual change detection to update element locators automatically.
Learning from Failures
Moving beyond learning exclusively from successes, future systems should systematically analyze failed attempts to refine skills proactively. When a concrete implementation fails, the agent could compare it against successful implementations of the same abstract skill to identify divergence points. Alternatively, it could request targeted human feedback on the specific failure mode rather than the entire skill, making human-in-the-loop learning efficient and practical.
Training Autonomous Skill Learners
Rather than relying on prompting large proprietary models, future research should investigate training smaller, open-source models to acquire polymorphic skills autonomously. RL offers a natural paradigm: agents could interact with web environments and receive rewards for discovering useful, generalizable skills. Key challenges include designing reward functions that encourage abstraction over specialization, solving temporal credit assignment when skill creation and reuse are separated by many steps, and developing efficient exploration strategies. Recent advances in hierarchical RL and continual learning provide promising starting points for this direction.
Collaborative Skill Ecosystems
Looking beyond individual agents, future systems could enable skill sharing across multiple agents through centralized or federated skill libraries. This raises interesting questions about skill quality control—how to validate contributions from different agents—version management for evolving skills, and personalization mechanisms that balance standardized, broadly applicable skills with agent-specific adaptations. Such collaborative approaches could dramatically accelerate learning across agent populations while maintaining the benefits of individual specialization.
7 Acknowledgment
We are grateful to our colleagues at Orby (now part of Uniphore) for their insightful discussions, comments, and support in setting up the evaluation environments during Simon’s internship. We also thank fellow Orby interns Yijin Ni and Mengzhao Jia for their valuable discussions and feedback. Additionally, we appreciate the members of the CHATS lab for their constructive feedback on an early version of this manuscript.
@@ 7 Acknowledgment
Appendix A Extended Related Works
@@
Skill Representation Formats The representation of a “skill” is a central design choice in skill-learning agents. The simplest forms are low-level and lack abstraction. Skills as Textual Descriptions leverage LLMs to store procedural knowledge in natural language within a prompt library (Wang et al., 2024; Zhu et al., 2025; Ouyang et al., 2025). This offers flexibility but lacks the formal structure needed for reliable composition or verification. Similarly, Skills as Action Traces store successful action sequences or workflows from prior tasks (Wang et al., 2024; 2025). While easy to record, these traces implicitly encode the specifics of a single website’s UI, making them fail when even minor elements change. A more structured approach represents Skills as Programs. This was pioneered in interactive environments like Minecraft, where agents build a library of reusable code snippets (Wang et al., 2023), and has been applied to web agents (Zheng et al., 2025) and software engineering (Chen et al., 2025a). However, these learned programs are typically concrete implementations tied to a single context, not inherently designed to handle variations across different websites fulfilling the same function.
@@
Appendix B Skill Examples
@@
Appendix C Experiment Details
C.1 Formal Definitions of Evaluation Metrics
@@ C.3 Dataset
C.4 Real-World Website Tasks
@@ C.5 Prompts
Appendix D Complete Results
@@
PolySkill (+Online) $\textbf{63.2}_{\pm 2.4}$ (+9.4) $\textbf{61.3}_{\pm 2.2}$ (+5.1) $\textbf{63.4}_{\pm 2.7}$ (+3.2) $\textbf{64.6}_{\pm 2.3}$ (+5.5) $\textbf{66.2}_{\pm 1.6}$ (+1.8) $\textbf{68.3}_{\pm 1.8}$ (+2.1)
@@
PolySkill (+Online Update) $\textbf{47.5}_{\pm 4.3}$ (+6.7) $\textbf{41.8}_{\pm 4.6}$ (+3.7) $\textbf{39.9}_{\pm 4.2}$ (+2.4) $\textbf{49.2}_{\pm 4.7}$ (+4.4) $\textbf{47.1}_{\pm 4.4}$ (+2.9) $\textbf{46.0}_{\pm 4.8}$ (+2.3)
@@ 3. Self-guided Exploration
Appendix E Extended Methodology
Algorithm 1 PolySkill: Polymorphic-Driven Skill Induction
3:for $t=1,\dots,N$ do
4:  Let $q_{t}$ be the current task from $\mathcal{Q}$.
9:   $\mathcal{A}_{t}\leftarrow\mathcal{A}_{t-1}\cup\mathcal{K}_{new}$ $\triangleright$ Update skill library for the next task
10:  else
11:   $\mathcal{A}_{t}\leftarrow\mathcal{A}_{t-1}$ $\triangleright$ Skill library remains unchanged on failure
12:  end if
13:end for
14:return $\mathcal{K}_{N}$

### v2

@@ 4.2 Analysis
To investigate how skills are induced during training, we use our newly proposed metrics to track the model’s learning dynamics. This analysis builds on previous work that sought to understand these dynamics (Shah et al., 2025). For our study, we saved 20 model snapshots at regular intervals while training on the Mind2Web Cross-task and WebArena Shopping benchmarks. We then evaluated our metrics on each snapshot to observe how skills emerge over time.
@@
To validate the fundamental hypothesis that skill learning improves task efficiency, we analyze the relationship between Skill Reusability and the number of steps required for task completion. The results, presented in Figure 4, demonstrate a clear inverse correlation across all three methods: as skill reusability increases from 0% to over 20%, the average number of steps decreases substantially from approximately 6.1 to 3.3-4.4 steps. PolySkill achieves the highest Skill Reusability (reaching 20.4% by task 180), while maintaining competitive step reduction. Notably, ASI shows the most dramatic step reduction despite lower utilization rates, suggesting efficient skill application. This analysis confirms that learned skills directly translate to improved task efficiency, with higher utilization rates consistently leading to more streamlined task execution. The strong correlation validates our core premise that skill induction and reuse are key drivers of agent performance improvement in complex web navigation tasks.
Number of Skills Learned
We additional analyze the size of the induced skill library (“# Skill”). Ideally, an agent should identify a concise set of “polymorphic” skills that can be reused across various contexts, rather than memorizing a vast library of specific, non-transferable sub-routines. A lower skill count coupled with high accuracy indicates that the agent has successfully distilled the underlying logic of the tasks in certain layouts. We measure the number of skills by measuring how many unique skills are set aside, without considering the exact implementation per class. We include both the w/o Update and w/ Update settings.
As shown in Table 2, PolySkill demonstrates superior efficiency compared to ASI. In the static setting, PolySkill achieves higher accuracy (55.4% vs. 52.3% in Cross-task) while maintaining a smaller skill library (43 vs. 50 skills). This confirms our hypothesis that polymorphic skills are more robust to interface variations, allowing a single skill to cover scenarios that would require multiple distinct skills in ASI.
Method Cross-task Cross-Website Cross-Domain
Acc $\uparrow$ # Skill $\downarrow$ Acc $\uparrow$ # Skill $\downarrow$ Acc $\uparrow$ # Skill $\downarrow$
Baseline 53.8 - 56.2 - 62.3 -
Static Approaches
ASI 52.3 50 54.9 47 57.3 33
PolySkill 55.4 43 57.6 44 60.1 36
w/ Online Update
ASI + Update 59.4 66 58.7 71 62.1 66
PolySkill + Update 63.2 47 61.3 53 63.4 56
Table 2: Performance comparison on Mind2Web using GPT-4.1. We report success rate (Acc $\uparrow$) and # Skill $\downarrow$.
When online updates are enabled, the skill library naturally grows to accommodate novel environments; however, PolySkill + Update continues to outperform the baseline significantly. Notably, in the Cross-task setting, PolySkill + Update reaches state-of-the-art performance (63.2%) with only 47 skills, far fewer than the 66 skills required by ASI + Update. This suggests that PolySkill adapts by refining existing polymorphic definitions rather than blindly accumulating redundant skills.
@@
Case Study: Continual Learning To simulate how skill would be used and evolved in real-world scenarios, we investigate the agent’s continual learning capabilities. The experiment begins with an agent whose skill library is initialized on the WebArena Shopping tasks. The agent then continues to perform online update via skill induction on new websites for Amazon and then Target, respectively (details in Section E.4). We performed the experiments 3 times to reduce the potential variance across run. In this setting, we are both interested in the positive transfer from the existing skill library in a similar domain. Also, we measure the WA Shopping performance after it has adapted to new website. This allows us to study potential catastrophic forgetting, where learning new knowledge can harm existing skills, a well-known challenge in continual learning (van de Ven et al., 2025).
@@ 5 Conclusion
Looking ahead, this framework opens several avenues for research, including handling highly dynamic websites, enabling skill sharing between agents, and incorporating human feedback. More broadly, the principle of polymorphism is not limited to web agents. It offers a powerful template for any agent that must operate in diverse environments sharing similar underlying structures (Xu et al., 2025)—from robotics, where skills must generalize across physical settings (Cheng et al., 2025; Liu et al., 2025a), to tool use, where software interfaces constantly evolve (Fei et al., 2025; Qiu et al., 2025). This work, therefore, provides a concrete step toward more robust and adaptive agents capable of learning from experience (Silver and Sutton, 2025) in diverse environments.
@@ 5.1 Future Work
To address the challenge of domains where task category boundaries are “fuzzy”, a prime candidate for future work is moving from manual definition to autonomous skill clustering. Instead of relying on predefined labels, agents could propose abstract classes based on functional similarity, i.e. via soft-clustering of execution traces, allowing abstract interfaces to emerge organically. Furthermore, we envision enabling smaller models to acquire these polymorphic skills autonomously through RL . In this paradigm, the efficiency penalty ($\gamma$) would evolve from a fixed heuristic into a mechanism for reward shaping, allowing the agent to dynamically learn the optimal trade-off between exploration effort and skill reusability.
6 Acknowledgement
We are grateful to our colleagues at Orby (now part of Uniphore) for their insightful discussions, comments, and support in setting up the evaluation environments during Simon’s internship. We also thank fellow Orby interns Yijin Ni and Mengzhao Jia for their valuable discussions and feedback. Additionally, we appreciate the constructive feedback from the members of the CHATS lab on an early version of this manuscript. We thank Modal Lab and Thinking Machines Lab for their generous sponsorship of computing resources.
Reproducibility Statement
The experiments on open-source models are run on an 8-H100 GPU cluster, and all calls for proprietary LLMs are made via their official API. All experiments are done via BrowserGym (de Chezelles et al., 2025) API, and experiment details are illustrated in Section 4 and Appendix E.
@@ 6 Acknowledgement
Appendix A Large Language Model Usage
We used large language models (LLMs) only for language refinement tasks, including grammar checking, phrasing adjustments, and enhancing readability. All scientific ideas, experiments, analyses, and results are the sole contributions of the authors. We only used LLMs for literature searching. It is important to note that the LLM was not involved in the ideation, research methodology, or experimental design. All research concepts, ideas, and analyses were developed and conducted by the authors.
Appendix B Extended Related Works
@@
Skill Representation Formats The representation of a “skill” is a central design choice in skill-learning agents. The simplest forms are low-level and lack abstraction. Skills as Textual Descriptions leverage LLMs to store procedural knowledge in natural language within a prompt library (Wang et al., 2024; Zhu et al., 2025). This offers flexibility but lacks the formal structure needed for reliable composition or verification. Similarly, Skills as Action Traces store successful action sequences or workflows from prior tasks (Wang et al., 2024; 2025). While easy to record, these traces implicitly encode the specifics of a single website’s UI, making them fail when even minor elements change. A more structured approach represents Skills as Programs. This was pioneered in interactive environments like Minecraft, where agents build a library of reusable code snippets (Wang et al., 2023), and has been applied to web agents (Zheng et al., 2025) and software engineering (Chen et al., 2025a). However, these learned programs are typically concrete implementations tied to a single context, not inherently designed to handle variations across different websites fulfilling the same function.
@@
Appendix C Limitations
We show that PolySkill also works for the latest agentic models for continual learning, which is a way to continuously improve these models with environment interaction settings. This also directly answers the claim that open-source models still fall behind in terms of skill induction (Shah et al., 2025). The promising next step would be to enable smaller agentic models to also utilize such skill acquisition behaviors, via further RL training on these memory environments.
Appendix D Skill Examples
@@
Appendix E Experiment Details
E.1 Formal Definitions of Evaluation Metrics
@@ E.3 Dataset
E.4 Real-World Website Tasks
@@ E.5 Prompts
Appendix F Complete Results
@@
PolySkill (+Update) $\textbf{63.2}_{\pm 2.4}$ (+9.4) $\textbf{61.3}_{\pm 2.2}$ (+5.1) $\textbf{63.4}_{\pm 2.7}$ (+3.2) $\textbf{64.6}_{\pm 2.3}$ (+5.5) $\textbf{66.2}_{\pm 1.6}$ (+1.8) $\textbf{68.3}_{\pm 1.8}$ (+2.1)
@@
PolySkill (+Update) $\textbf{47.5}_{\pm 4.3}$ (+6.7) $\textbf{41.8}_{\pm 4.6}$ (+3.7) $\textbf{39.9}_{\pm 4.2}$ (+2.4) $\textbf{49.2}_{\pm 4.7}$ (+4.4) $\textbf{47.1}_{\pm 4.4}$ (+2.9) $\textbf{46.0}_{\pm 4.8}$ (+2.3)
@@ 3. Self-guided Exploration
F.1 Cost and Efficiency Analysis
To evaluate the economic feasibility and efficiency of our framework, we conducted a detailed cost analysis measuring token consumption. We calculated the average token usage (input + output) per instance in the Mind2Web cross-task setting using GPT-4.1. The breakdown distinguishes between the Action Generation phase (inference) and the overhead costs associated with Verification and Skill Induction.
Table 10: Average token consumption per task instance on Mind2Web. PolySkill significantly reduces token usage during action generation, leading to lower long-term costs despite the initial overhead of skill induction.
Method Action Gen. Verification Skill Induction Total Tokens
Baseline (No Skills) 55,367.4 55,367.4
ASI 53,840.5 4,673.8 5,773.2 64,287.5
PolySkill 49,710.3 4,963.2 8,476.4 63,149.9
As shown in Table 10, PolySkill introduces an overhead for skill induction and verification, resulting in a higher total token count for the initial run compared to the no-skill baseline. However, this investment yields significant efficiency gains:
  1. 1.
    Reduced Inference Cost: PolySkill reduces token consumption during the Action Generation phase by approximately 10.2% compared to the Baseline (from $\sim$55.4k to $\sim$49.7k tokens) and 7.7% compared to ASI. This is attributed to the polymorphic skills allowing the agent to solve tasks in fewer steps with higher precision.
  2. 2.
    Amortized Efficiency: Crucially, the costs for Verification and Skill Induction are primarily one-time (or amortized) costs. Once a concrete implementation for a website is induced, it is stored in the library and reused for all future tasks on that site. Consequently, as the number of tasks performed on a website increases, the average cost per task converges toward the significantly lower Action Generation cost, making PolySkill highly cost-effective for long-term deployment.
Appendix G Extended Methodology